OpenStudy (anonymous):

Triangle PQR is a right triangle. If RQ=8, what is PQ? A. 4 radical 3 B. 16 C. 4 D. 8 radical 3

OpenStudy (anonymous):



OpenStudy (anonymous):

Using cos (angle) = Adjacent / Hypotenuse cos 60 = RQ / PQ cos 60 = 8 / PQ PQ = ?????

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Lol Yw.

OpenStudy (anonymous):

sin(30) = RQ/PQ sin(30) = 8/PQ 1/2 = 8/PQ PQ = 2*8 PQ = 16